OpenStudy (anonymous):

−4(x + 1) − 3 = −3(x − 4)

OpenStudy (tkhunny):

Use your distributive property - once on each side.

OpenStudy (anonymous):

After distributing you will get -4x - 4 - 3 = -3x + 12 - 4x - 7 = -3x + 12 - 4x = - 3x + 19 -1x = 19 x = - 19

OpenStudy (anonymous):

-4x + -4 - 3 + -3x - (-12)